Grade 12th passPhysical Chemistry

3 moles of A and 4 moles of B are mixed together and allowed to come into equilibrium according to the following reaction.
3A(g) + 4B(g) = 2C(g) +3D(g)
when equilibrium is reached, There is 1 mole of C. The equilibrium constant of the reaction is
  1. 1÷4
  2. 1÷3
  3. 1÷2
  4. 1
​Please explain step by step...............

Yasir Rather
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
3A + 4 B = 2C + 3 D
 
at equilibrium
3- 3x  4- 4x    2x     3x
 
as 2x = 1
x = ½
 
now equilibrium constant = 6x^2 / (3- 3x) (4-4x) = 6/4 / 3 = 2/4 = ½
 
hence option C is correct